OpenStudy (anonymous):

whats the derivative of y=x-2cosx

OpenStudy (anonymous):

Do you know the derivative of \(x\)? What about \(2\cos x\)?

OpenStudy (anonymous):

You just need to distribute the differential. so it's the derivative of x, and the derivative of -2cos(x). the derivative of x is 1, and the derivative of -2cos(x) is 2sin(x)

OpenStudy (anonymous):

final answer should be 1 + 2sin(x)

OpenStudy (anonymous):

just remember that you are taking the derivative in respect to x because that is the only variable of the function derivative of x is 1 and derivative of -2cos(x) is 2sin(x) put it together and you have \[f \prime=1+2\cos(x)\]